The cycle work of an air standard engine is 1000kJ/kg.The maximum cycle temperature is 3200 degree C and the temperature at the end of isentropic compression is 650 degrees Celsius. Calculate the compression ration of the engine
1
Expert's answer
2021-07-19T02:29:01-0400

We know that

"PV^{\gamma-1}=constant"

"(\frac{V_2}{V_1})=\frac{T_2}{T_2}^{\gamma-1}"

Compressions ratio (r)="\frac{V_2}{V_1}"


"r=(\frac{T_2}{T_1})^{\gamma-1}=(\frac{3473}{923})^{1.5}=3.76^{1.5}=7.2"